4x +4y+8xy+3=0.Which conic does this equation represent?

Grade:12th pass

1 Answers

Harsh Patodia IIT Roorkee
6 years ago
here a=0,b=0,g=2,f=2,h=4,c=3 from comparing with two degree equation
if abc+2hfg-af2-bg2-ch2=0 represent a pair of lines otherwise a conic
here 2(4)(2)(2)-3(4)(4) not equal to zero so its a conic now here h2-ab=16>0 so its hyperbola (a+b=0 its a rectangular hyperbola)